# Unveiling Parts Beyond Objects: Towards Finer-Granularity Referring Expression Segmentation

### 13 Dec 2023 v1

@@ 1 Introduction
Since the concept of RES task was initially proposed in [13], various multimodal frameworks such as [13, 41, 43, 19] towards precise RES have been proposed to deal with the challenging feature extraction and alignment problems between visual and linguistic modalities. However, current works are limited to the scope of object-level visual grounding. As shown in Table 1, the RefCOCO dataset [45] stands as one of the most widely used grounding benchmarks by far, basically containing only object-level visual and textual annotations. It does not take into account the part-level grounding task, which is crucial for future multimodal models to act as intelligent agents to realize the fine-grained perception of the real world. The existing visual grounding datasets couldn’t support the training and evaluation of such advanced capabilities. Thus, exploring how to transcend the current object-level visual grounding constraints and delve into finer-grained part-level grounding, is highly meaningful and worthy of in-depth study, which is the primary focus of this work.
@@ 4.2 MRES-32M Dataset Details
Figure 4: Selected samples of our proposed MRES-32M dataset.
As listed in Table 2, existing datasets, such as the most commonly used benchmark dataset RefCOCO [45], have limitations in terms of small data scale and lacking part-level dense annotations. As one of the pioneering works for finer-granularity unimodal visual comprehension, PACO [33] only contains semantic tags of objects or parts, without informative descriptions and visual context encompassed as referring expressions. Summarily, we compare the proposed MRES-32M with existing datasets and list some unique and significant properties of our dataset in Table 2. Besides, we have also provided a few examples in our MRES-32M dataset in Fig. 4.
@@
Breakable Data Scales. To the best of our knowledge, MRES-32M is the largest dataset in the current grounding research community. In terms of the number of images, object instances and referential tokens, it surpasses the largest existing visual grounding dataset RefCOCOg [27] by factors of 50, 158, and 184, respectively. Meanwhile, it encompasses part instance counts that exceed the largest existing part semantic segmentation dataset [33] by 58 times.
@@ 6 Experimental Results
To evaluate the effectiveness and the designing rationale of our method, comprehensive experiments are conducted on three datasets (i.e., RefCOCO [45], RefCOCO+ [45], RefCOCOg [27]) for classic RES and our newly built RefCOCOm benchmark dataset for multi-granularity RES.
@@ 6.2 Experimental Setup
UniRES (Ours) 79.2 81.6 76.6 73.0 78.1 65.8 71.8 72.9
6.3 Main Results
6.3.1 Classic Object-Level RES Task
Comparison with SOTA Methods. To validate the superiority of the constructed MRES-32M dataset and our model UniRES, our framework is fairly evaluated against previous SOTA methods on RefCOCO [45], RefCOCO+ [45] and RefCOCOg [27] datasets under the settings of both zero-shot and fine-tuning. As presented in Table 3, our method greatly outperforms previous methods in terms of segmentation accuracy across all evaluation subsets of the three benchmark datasets. Via directly zero-shot transferring to the downstream classic RES task after pre-training on our MRES-32M dataset, Our UniRES model achieves a leading zero-shot segmentation accuracy (i.e., approximately 71%) without any fine-tuning to adapt to downstream task’s data, which is significantly higher than all the recently proposed zero-shot RES methods by a large margin (i.e., ↑30-40% mIoU). At the same time, it is worth noting that our zero-shot segmentation performance is already better than many of previous RES methods under fine-tuning setting (e.g., CRIS [41] and ReSTR [19]), validating the potential of our MRES-32M dataset and UniRES model. Furthermore, after fine-tuning on the specific downstream datasets for classic RES task, our UniRES greatly surpasses all the previous RES methods no matter the specialist models (e.g., LAVT [43] and LISA [22]) for classic RES task or the powerful generalist models (e.g., X-Decoder [49] and SEEM [50]).
6.3.2 Multi-Granularity MRES Task
Methods val testA testB
Part Obj & Part Part Obj & Part Part Obj & Part
Specialists
SeqTR [48] 13.9 28.2 12.1 22.8 18.1 34.7
CRIS [41] 10.6 25.4 10.1 21.2 12.9 30.0
LAVT [43] 15.3 29.9 13.2 24.4 18.7 35.5
Generalists
X-Decoder [49] 16.2 29.5 13.6 23.6 20.3 33.8
SEEM [50] 16.1 29.4 13.6 23.4 20.4 33.9
UniRES (Ours) 19.6 34.3 16.4 27.8 25.2 41.7
@@ 6.4 Ablation Studies
We conduct extensive ablation experiments on the validation set of our RefCOCOm benchmark in terms of mIoU. The tables below involves three evaluation settings on our RefCOCOm, including using only objects as RES targets, using only parts as RES targets, and a mixed setting that combines both of the above two granularity levels.
@@ A.5 More Visualization Results
Samples of MRES-32M Dataset. A few more examples in our newly built MRES-32M dataset for visual grounding task are presented in Fig. 10.
Figure 10: More selected samples from our built MRES-32M dataset for visual grounding tasks.

### 21 Mar 2024 v2

@@ 1 Introduction
Since the concept of RES task was initially proposed in [13], various multimodal frameworks such as [13, 41, 43, 19] towards precise RES have been proposed to deal with the challenging feature extraction and alignment problems between visual and linguistic modalities. However, current works are limited to the scope of object-level grounding. As shown in Table 1, the RefCOCO dataset [45] stands as one of the most widely used grounding benchmarks by far, basically containing only object-level visual and textual annotations. It does not take into account the part-level grounding task, which is crucial for future multimodal models to act as intelligent agents to realize the fine-grained perception of real world. The existing grounding datasets couldn’t support the training and evaluation of such advanced capabilities. Thus, exploring how to transcend the current object-level constraints and delve into finer-grained part-level grounding, is highly meaningful and worthy of in-depth study, which is the primary focus of this work.
@@ 4.2 MRES-32M Dataset Details
As listed in Table 2, existing datasets, such as the most commonly used benchmark dataset RefCOCO [45], have limitations in terms of small data scale and lacking part-level dense annotations. As one of the pioneering works for finer-granularity unimodal visual comprehension, PACO [33] only contains semantic tags of objects or parts, without informative descriptions and visual context encompassed as referring expressions. Summarily, we compare the proposed MRES-32M with existing datasets and list some unique and significant properties of our dataset in Table 2. Besides, we have also provided a few examples in our MRES-32M dataset in the appendix.
@@
Breakable Data Scales. To the best of our knowledge, MRES-32M is the largest dataset in the current grounding research community. In terms of the number of images and object instances, it surpasses the largest existing visual grounding dataset RefCOCOg [27] by factors of 38 and 283, respectively. Meanwhile, it encompasses part instance counts that exceed the largest existing part semantic segmentation dataset [33] by 58 times.
@@ 6 Experimental Results
To evaluate the effectiveness of our method, comprehensive experiments are conducted on three classic RES datasets (i.e., RefCOCO [45], RefCOCO+ [45], RefCOCOg [27]) and our RefCOCOm benchmark for multi-granularity RES.
@@ 6.2 Experimental Setup
6.3 Main Results
6.3.1 Multi-Granularity MRES Task
Methods val testA testB
Part Obj & Part Part Obj & Part Part Obj & Part
Specialists
SeqTR [48] 13.9 28.2 12.1 22.8 18.1 34.7
CRIS [41] 10.6 25.4 10.1 21.2 12.9 30.0
LAVT [43] 15.3 29.9 13.2 24.4 18.7 35.5
Generalists
X-Decoder [49] 16.2 29.5 13.6 23.6 20.3 33.8
SEEM [50] 16.1 29.4 13.6 23.4 20.4 33.9
UniRES (Ours) 19.6 34.3 16.4 27.8 25.2 41.7
6.3.2 Classic Object-Level RES Task
@@
UniRES (Ours) 79.2 81.6 76.6 73.0 78.1 65.8 71.7 73.2
Comparison with SOTA Methods. To validate the superiority of our MRES-32M dataset and model UniRES, our framework is fairly evaluated against previous SOTA methods on RefCOCO [45], RefCOCO+ [45] and RefCOCOg [27] under both zero-shot and fine-tuning settings. As presented in Table 4, our method greatly outperforms previous methods in terms of segmentation accuracy across all the benchmark datasets. Via directly zero-shot transferring to the downstream classic RES task after pre-training on our MRES-32M dataset, Our UniRES model achieves a leading zero-shot segmentation accuracy (i.e., approximately 71%) without any fine-tuning to adapt to downstream task’s data, which is significantly higher than all the recently proposed zero-shot RES methods by a large margin (i.e., ↑30-40% mIoU). At the same time, it is worth noting that our zero-shot segmentation performance is already better than many of previous RES methods under fine-tuning setting (e.g., CRIS [41] and ReSTR [19]), validating the potential of our MRES-32M dataset and UniRES model. Furthermore, after fine-tuning on the classic RES datasets, our UniRES greatly surpasses previous methods no matter the specialist models (e.g., LAVT [43] and LISA [22]) for classic RES or the generalist models (e.g., X-Decoder [49] and SEEM [50]).
@@ 6.4 Ablation Studies
We conduct ablation experiments on our RefCOCOm validation set. The tables below involves three evaluation settings on our RefCOCOm, using only objects as RES targets, using only parts as RES targets, and a mixed setting that combines both of the above two granularity levels.
@@ 7 Conclusion and Broader Impact
Acknowledgement
We thank Yepeng Tang for the helpful discussions on this work, all the Image & Video Analysis Group (IVA)’s members in CASIA for the technical support, and all the insightful reviewers for the helpful suggestions. This work was supported by the National Science and Technology Major Project (No.2022ZD0118801), National Natural Science Foundation of China (U21B2043, 62206279).
@@ A.5 More Visualization Results
Samples of MRES-32M Dataset. A few examples in our newly built MRES-32M dataset for visual grounding task are presented in Fig. 9.
Figure 9: Selected samples from our built MRES-32M dataset for visual grounding tasks.